A solution was prepared by mixing 10.0 g of NaCl in 100.0 g of water. What is the percent concentration of the solution?
Answer #1


mass(NaCl)= 10.0 g


mass(solvent)= 100.0 g

mass of solution = 10.0+100.0 = 110.0 g

use:
mass% = mass of NaCl *100 / mass of solution
= 10.0*100/110.0
= 9.091 %
Answer: 9.09 %